Walt made an extra $9000.00 last year from a part time job. He invested part of the money at 8% and the rest at 7%. He made a total of $680.00 in nterest. How much was invested?
Answer by mananth(16946)

let him invest $x at 8%
he will invest 9000-x at 7%
he recieved an interest of $680
..
0.08x+0.07(9000-x)=680
multiply by 100
8x+7(9000-x)=100*680
8x+63000-7x= 68000
x= 68000-63000
x= 5000 amount invested @ 8%
balance $4000 he invests at 7%